OpenStudy (anonymous):

The two-column proof with missing justifications proves the Pythagorean Theorem using similar triangles.

OpenStudy (anonymous):

Statement Justification Draw an altitude from point C to Line Segment AB Let Line Segment BC= a Line Segment CA= b Line Segment AB= c Line Segment CD= h Line Segment DB= x Line Segment AD= y y + x = c c over a is equal to a over y; c over b is equal to b over x a2 = cy; b2 = cx a2 + b2 = cy + b2 a2 + b2 = cy + cx a2 + b2 = c(y + x) a2 + b2 = c(c) a2 + b2 = c2 Which is NOT a justification for the proof? Addition Property of Equality Pythagorean Theorem Pieces of Right Triangles Similarity Theorem Cross Product Property

OpenStudy (anonymous):

Given: ∆BCA is a right triangle. Prove: a2 + b2 = c2

OpenStudy (anonymous):

Will give medal out to best answer

OpenStudy (anonymous):

@jim_thompson5910